OpenStudy (anonymous):

For the quadratic function, finds it's vertex and line of symmetry, then graph the function. 3x^2 - 18x +32.

OpenStudy (amistre64):

it vertex will be on its line of symmetry

OpenStudy (amistre64):

given a quad of the form: ax^2 + bx + c the line of symmetry can be formulated as: x = -b/2a by plugging value of x into the eqaution, we can solve for y

OpenStudy (anonymous):

Ok I am stuck at this \[\frac{ 18 \pm \sqrt{-60} }{ 6}\] I know it will be \[3 \pm \sqrt{-60}\] but do I change that -60 to \[2\sqrt{5i}\] over 6?

OpenStudy (anonymous):

Ok I guess what I'm asking is my vertex going to be \[3 + \frac{ 2\sqrt{5i} }{ 6 }, 3 - \frac{ 2\sqrt{5i} }{ 6 }\]

OpenStudy (amistre64):

there are 2 parts to the formula you are trying to use. there is the axis of symmetry part and the distance from the axis